# OnePerc: A Randomness-aware Compiler for Photonic Quantum Computing

### v2

@@ 3. Overview
To provide a concise overview, we exemplify the compilation flow by compiling the program graph state in Fig. 4(a) onto the hardware in Fig. 4(b), which is 3 layers of that in Fig. 3(b). Indeed, while Fig. 3(b) depicts only a single RSL, the incorporation of additional layers is both allowed and necessary for larger graph states. Steps (b)$\rightarrow$(d)$\rightarrow$(c) demonstrate the online pass, while step (a)$\rightarrow$(c) illustrates the offline pass. In the online pass, fusions are conducted concurrently in a predetermined pattern (Fig. 4(d)) without individual retries of the failed ones, which eliminates the necessity for sequential operations. This results in a physical graph state with a random 2D structure, which is then reshaped to a program-agnostic $2\times 1$ lattice (Fig. 3(c)), characterized by a regular structure to enhance real-time efficiency. When the fusion success probability exceeds the percolation threshold, this reshaping process attains near-deterministic success as the RSL size increases. This eliminates the necessity for repetitive retries of the entire compilation. With this determinism, the offline pass can be employed to improve the efficiency by mapping the program graph state compactly onto the $2\times 1$ lattice (bold blue lines in Fig. 3(c)).
Note that the compilation of general programs can be considerably more intricate than the example presented here. First, the fusion strategy among resource states is more complex than Fig. 4(d). Specifically, it enables the formation of a 3D structure rather than 2D, being adaptable to various resource states and allowing pipelined feed-forward with a small overhead. Second, the complexity of the reshaping algorithm is carefully reduced to enhance its real-time scalability. This is achieved by improving the parallelism with a modular design on each RSL. Third, the reshaping process is heterogeneous between the spatial and temporal dimensions, with the temporal dimension supporting connections both between adjacent layers or non-adjacent layers. These flexible connections provides a larger optimization space for the offline mapping than Fig. 4(c). Forth, the online and offline passes are further bridged by posing a FlexLattice IR, which guides the low-level operations by its translation to an instruction set. For general programs, the compilation flow can be summarized as follows.
@@
    Before program execution begins, an offline pass maps the program graph state to an efficient FlexLattice IR program, which translates to low-level instructions to guide real-time operations (Section 6).
@@
    During real-time execution, fusions between resource states are performed concurrently in a predetermined pattern that allows pipelined feed-forwards to improve the long-range-connectivity of the resulting graph state (Section 4).
@@
    The resulting random physical graph state is then reshaped to a 3D structure to fulfill the requirement of the IR program, with measurements performed on qubits according to the IR program (Section 5).
@@ 4. Resource State Fusion
From the motivating example, it becomes evident that a dynamic adjustment of the fusion strategy among resource states lacks real-time scalability, as its prolonged decision making time could easily exceed the photon lifetime. To address this, we opt for a semi-static fusion strategy that generates program-agnostic physical graph states. This strategy is static in that it is predetermined independent of high-level programs, yet semi-static in that it allows pipelined feed-forward which induces only a constant overhead.
@@ 4.1. Sufficient / Insufficient Degree
The predetermined strategy attempts to create a lattice structure from the resource states, which is straightforward when resource states have sufficient node degrees, i.e., the maximum degree in the resource states surpasses that in the lattice. Fig. 5(a) shows the strategy of forming a 2D square lattice using 4-degree resource states. On the left side, each light blue line denotes a leaf-leaf fusion (degree=1) of the resource states (yellow circles). The right side displays the resulting random graph state due to fusion failures, with the consequences of successful and failed fusions depicted in the middle box. Similarly, Fig. 5(b) demonstrates the case of forming a cubic lattice 6-degree resource states. As resource states on realistic hardware may lack sufficient degrees (fbqc, ), such as forming 3D cubic lattices using 4-degree resource states, we can increase resource state degrees by merging multiple RSLs into one layer using root-leaf fusions between resource states, represented by the dark blue lines in Fig. 5(c). Upon fusion success, a root-leaf fusion between two 4-degree resource states can generate a 7-degree resource state.
@@ 4.2. Removal of Irregular Structures
However, a failed root-leaf fusion may result in irregular cyclic structures in the generated graph state, leading to significant challenges for subsequent reshaping process. For example, a root-leaf fusion between two resource states $A_{0}$ and $B_{0}$ in Fig. 6 will generate a tree-like resource states $A$ and a fully-connected cyclic $B$, which may result in complex structures if other qubits in $B_{0}$ were fused with other resource states. To remove these cyclic structures, we will show that resource state $B$ can be transformed to its local complementation, i.e., a tree-like resource state $C$ by adjusting the bases of subsequent measurements and fusions. The operators for local complementation can therefore be propagated to the end of the computing process, and eventually can be removed by a re-interpretation of the final outcomes.
@@
Definition 4.1 (local complementation).
Letting $G=(V,E)$ be a graph and $v\in V$, the local complement of G at $v$, denoted by $\tau_{v}(G)$, is obtained by complementing (inversing) the edges in the subgraph of $G$ induced by the neighborhood $N_{v}$ of a and leaving the rest of the graph unchanged.
Proposition 4.2 ((LCRule, ) LC-rule).
A graph state $G$ can be transformed to its local complementation $\tau_{v}(G)$ by applying the following sequence of single-qubit Clifford operations
$U_{v}(G)=\exp\large(-\mathrm{i}\frac{\pi}{4}X_{v}\large)\prod_{u\in N_{v}}\exp\large(\mathrm{i}\frac{\pi}{4}Z_{v}\large)$
where $N_{v}$ is the set of neighbors of $v$ in $G$.
Theorem 4.3 ().
The operators $U_{Z}^{\pm}=\exp(\pm\mathrm{i}\frac{\pi}{4}Z)$ and $U_{X}^{\pm}=\exp(\pm\mathrm{i}\frac{\pi}{4}X)$ can both be propagated through a single-qubit equatorial measurement on the Bloch sphere, i.e., a measurement in the basis of $\cos\phi X+\sin\phi Y$ where $\phi\in[0,2\pi)$, as well as a $Z-$measurement by a change of measurement basis.
@@
A measurement of $A$ on a single-qubit state $\left|\psi\right\rangle$ collapses it to a state $\left|\psi^{\prime}\right\rangle\equiv\mathrm{M}_{[A]}\left|\psi\right\rangle=\frac{\mathbb{I}\pm A}{2}\left|\psi\right\rangle$ when the outcome is 0 and 1, respectively. It can then be verified that
@@
$\mathrm{M}_{[\cos\phi X+\sin\phi Y]}U_{X}^{\pm}$ $=U_{X}^{\pm}\mathrm{M}_{[\cos\phi X\pm\sin\phi Z]}$
$\mathrm{M}_{Z}U_{Z}^{\pm}$ $=U_{Z}^{\pm}\mathrm{M}_{Z}$
$\mathrm{M}_{Z}U_{X}^{\pm}$ $=U_{X}^{\pm}\mathrm{M}_{Y}\qed$
Theorem 4.4 ().
The operators $U_{Z}^{\pm}=\exp(\pm\mathrm{i}\frac{\pi}{4}Z)$ and $U_{X}^{\pm}=\exp(\pm\mathrm{i}\frac{\pi}{4}X)$ can both be propagated through a two-qubit $XZ,ZX$ fusion, i.e., a joint measurement of $X_{1}Z_{2},Z_{1}X_{2}$ on qubit 1 and qubit 2, by a change of fusion basis. In particular, given a two-qubit quantum state $\left|\psi\right\rangle$,
@@
A measurement of $A$ on a two-qubit state $\left|\psi\right\rangle$ collapses it to a state $\left|\psi^{\prime}\right\rangle\equiv\mathrm{M}_{[A_{1}A_{2}]}\left|\psi\right\rangle=\frac{\mathbb{I}\pm A_{1}A_{2}}{2}\left|\psi\right\rangle$ when the outcome is 0 and 1, respectively. It can then be verified that
@@
$\mathrm{M}_{[X_{1}Z_{2}]}$ $\mathrm{M}_{[Z_{1}X_{2}]}U_{X_{1}}^{\pm_{1}}U_{X_{2}}^{\pm_{2}}=U_{X_{1}}^{\pm_{1}}U_{X_{2}}^{\pm_{2}}\mathrm{M}_{[\pm_{1}Y_{1}X_{2}]}\mathrm{M}_{[\pm_{2}X_{1}Y_{2}]}$
@@
4.3. Pipelined Feed-forward
Despite being predetermined, the semi-static fusion strategy allows collective feed-forwards which can be pipelined to reduce the overhead. On one hand, this is because the propagation of operators for local complementation requires an adaptive adjustment of measurement and fusion bases. One the other hand, the connectivity of the generated random graph state can be enhanced by mitigating the fusion failures, including retrying failed leaf-leaf fusions with redundant degrees (e.g., the $7^{\mathrm{th}}$ degree of $A_{1}$ in Fig. 6) and retrying failed root-leaf fusions with remaining degrees (e.g., $A$ and $C$ in Fig. 6). These feed-forwards can be pipelined as they involve only one or few consecutive RSLs. For example, the second feed-forward round of a (merged) RSL can be conducted concurrently with the first round of the next one, thereby introducing only a constant overhead for the overall program execution.
@@ 5. Random State Reshaping
Figure 7. (2+1)-D renormalization for handling random graph states generated by fusions.
Similar to the impracticality of dynamically adjusting fusion strategies, dynamically mapping program graph states onto the random physical graph states demands computational complexity exceeding the photon lifetime. To address this, we reshape each RSL to a program-agnostic 2D lattice structure and preserve flexibility for the connections among them. This (2+1)-D design, motivated by the continuous generation of RSLs over time and the presence of delay lines, aims to minimize the dynamic feed-forward while providing enough room for program optimization.
@@ 5.1. Efficient 2D Renormalization
On each (merged) RSL, we apply a process known as renormalization, which reshapes the largest connected component of the random physical graph state to a coarse-grained 2D lattice. The key to its viability lies in the percolation phenomenon (3GHZ, ; perc_threshold, ; 2D_renorm, ). That is, when the fusion success (or edge connectivity) probability exceeds a certain threshold, the random physical graph state undergoes a phase transition from short-range connectivity to long-range connectivity, leading to the largest connected component reaching a comparable size with the original graph state. Since fusions on each RSL are constrained as a squared lattice, the percolation threshold is only 0.5 (square_lattice_perc, ), lower than the practical fusion success probability.
Identifying intersections of horizontal and vertical paths within this component reveals a coarse-grained square lattice, represented by bold nodes and edges in Fig. 7(b). Specifically, we search for vertical paths from left to right and horizontal paths from bottom to top, enforcing distinct vertical or horizontal paths to maintain a minimum separation of at least one qubit. For each vertical (horizontal) path, a connectivity check is conducted between nodes at the top (left) and bottom (right), facilitated by a disjoint-set data structure to reduce the complexity. Upon confirming connectivity, a breadth-first search (BFS) is applied to determine the shortest path, ensuring it remains free of self-tangling. To further prevent tangling between vertical and horizontal paths, we remove the surrounding qubits of each identified path after discovery, preventing their interference with subsequent searches. Considering the removals, an alternating search of vertical and horizontal paths emerges as an effectively searching order.
To improve real-time scalability, the 2D renormalization is designed to allow modularity, with areas on the RSL renormalized concurrently and then joined together. As shown in Fig. 8, the RSL is divided into several modules of size $L_{Module}\times L_{Module}$, with some intervals of length $L_{interval}$ left in between for joining the modules with connected paths. This modularity leads to a time complexity of $O(1)$ constrained by the size of modules, with the complexity within each module further reduced by employing a disjoint-set data structure. However, it also introduces a resource overhead due to the presence of intervals and the possible missing connected paths between the modules . A suitable ratio of $L_{Module}$ and $L_{Interval}$, defined as MI ratio, can help mitigate this resource overhead, as will be demonstrated Section 7.
@@
5.2. Flexible Time-like Connection
Nodes on the renormalized 2D lattices can be connected along the third dimension of time, referred to as time-like connections. These connections can be established both between adjacent 2D lattices or across non-adjacent 2D lattices, which are called adjacent-layer connections and cross-layer connections. The connections to establish are given by the IR program at the beginning of execution, leading to a 3D graph consisting of interconnected 2D layers as illustrated in Fig. 7(a). The process of generating this 3D graph is illustrated in Fig. 7(c) on 8 RSLs, with the adjacent-layer connection $AC$ and the cross-layer connection $BD$ being implemented through the bold purple paths.
In particular, this process involves an attempt of 2D renormalization on each RSL. The successful ones then serves as a logical layer, which are indexed with integers in Fig. 7(c), while the failed ones serve as routing layers, which are indexed with decimals in Fig. 7(c). The renormalization on an RSL is considered successful if:
@@ 5.2. Flexible Time-like Connection
    The renormalized 2D lattice reaches a size of choice, equivalent with a choice of average node size, where
    $\mathrm{average\_node\_size}=\frac{\mathrm{RSL\_size}}{\mathrm{renormalized\_lattice\_size}}$
@@
    The RSL can establish necessary time-like connections with prior logical layers.
To establish a time-like connection between two nodes, a set of qubits around the earlier node are fused with corresponding qubits on a correct subsequent layer. For adjacent-layer connection such as $AC$, the set of qubits around $A$ are directly fused to the next layer, which is layer 1.1 in Fig. 7(c). For cross-layer connection such as $BD$, the set of qubits around $B$ are temporarily stored in delay lines as represented by the green thin lines in Fig. 7(c), until they can be fused to layer 2.1, which is the first RSL between the layer of the other node $D$ (layer 3) and its prior logical layer (layer 2). In contrast, all qubits of every path routing layers are directly fused with their next layer, as depicted by the grey thin lines in Fig. 7(c).
@@
Subsequently, a path searching between the two nodes is conducted within the random physical graph state, exemplified by the bold purple lines $AC$ and $BD$ in Fig. 7(c). Again, this is achieved by a connectivity check utilizing a disjoint-set data structure structure and a BFS for the shortest path. If the connectivity check yields a negative result, it indicates that the current RSL fails to meet the second condition, thus the RSL would serve as a routing layer. It is worth mentioning that this process can tolerate the photon loss by treating a fusion as successful only if both the input photons are detected. Effectively, the presence of photon loss causes a reduction of the fusion success probability, possibly leading to more path routing layers between logical layers.
6. IR for Offline Optimization
The structures supported by the reshaping of the random physical graph state motivates a new IR, which maintains the high-level program information and allows sophisticated offline optimizations. Specifically, an offline mapping algorithm can be employed to transform the program graph state to an IR graph state. This IR will then guide the real-time operations through a low-level instruction set.
6.1. FlexLattice IR
The reshaped physical graph state enables the abstraction of a deterministic virtual hardware, characterized by 2D layers in a lattice structure that are flexibly interconnectable in the third dimension, as depicted in Fig. 9(b). This virtual hardware allows for the embedding of a family of graphs sharing similar features as demonstrated in Fig. 9(c), motivating the introduction of a FlexLattice IR. The purpose of this IR is to preserve the high-level program information while providing a maximal optimization space supported by the hardware. The rationale for this IR is supported by the fact that any program graph state can be transformed into such a representation, as illustrated in Fig. 9(a)$\rightarrow$(c)$\rightarrow$(d).

### v3

@@ 3.2. Framework Overview
Note that the compilation of general programs can be considerably more intricate than the example presented here. First, the fusion strategy among resource states is more complex than Fig. 6(d). Specifically, it enables the formation of a 3D structure rather than 2D, being adaptable to various resource states and allowing collective retries with a small overhead. Second, the complexity of the reshaping algorithm is carefully reduced to enhance its real-time scalability. This is achieved by a modular design on each RSL that improves the parallelism of path searching. Third, the reshaping process is heterogeneous in the spatial and temporal dimensions, with the temporal dimension supporting connections both between adjacent layers and non-adjacent layers. These flexible connections provides a larger optimization space for the offline mapping than Fig. 6(c). Forth, the online and offline passes are further bridged by posing a FlexLattice IR, which guides the low-level operations by its translation to an instruction set. For general programs, the compilation flow can be summarized as the following.
@@
    Before program execution, an offline pass transforms the program graph state to an efficient FlexLattice IR, which is then translated to intermediate-level instructions to guide real-time operations (Section 6).
@@
    During real-time execution, fusions between resource states are performed concurrently in a predetermined pattern, allowing collective retries of failed connections to improve the long-range-connectivity of the resulting physical graph state (Section 4).
@@
    The resulting physical graph state is then reshaped to a 3D structure that fulfills the requirement of the IR program, with measurements performed on qubits according to the IR program (Section 5).
The following sections (Section 4, 5, 6) will provide a bottom-up introduction to the framework.
@@ 4. Resource State Fusion
In this section, we discuss the semi-static fusion strategy for generic star-like resource states. This strategy is static in that it is predetermined independent of high-level programs, yet semi-static in that it allows collective retries which induces only a constant overhead. In Fig. 2, this corresponds to the online pass from resource states to physical graph states.
@@ 4.1. Sufficient / Insufficient Degree
The predetermined strategy attempts to create a lattice structure from the resource states, which is straightforward when resource states have sufficient node degrees, i.e., the maximum degree in the resource states surpasses that in the lattice. For example, Fig. 7(a) shows the strategy of forming a 2D square lattice using 4-degree resource states. On the left side, each light blue line represents a leaf-leaf fusion of the resource states (yellow circles). The right side displays the resulting physical graph state, with the consequences of successful and failed fusions depicted in the middle box. Similarly, Fig. 7(b) demonstrates the case of forming a 3D cubic lattice from 6-degree resource states.
As resource states on realistic hardware may lack sufficient degrees, such as forming 3D cubic lattices using 4-degree resource states, we can increase resource state degrees by merging multiple RSLs into one layer using root-leaf fusions, represented by the dark blue lines in Fig. 7(c). Upon fusion success, while the two qubits in the fusion vanish, the two set of neighboring qubits of them would be connected in pairwise. Hence a successful root-leaf fusion between two 4-degree resource states can generate a 7-degree graph state, which then has sufficient degree to form a 3D lattice.
@@ 4.2. Removal of Irregular Structures
However, a failed root-leaf fusion may result in irregular cyclic structures in the generated graph state, leading to significant challenges for subsequent reshaping process. For example, a failed root-leaf fusion between two resource states $A_{0}$ and $B_{0}$ in Fig. 8 generates a star-like graph state $A$ and a fully connected cyclic graph state $B$. This is because a failed fusion on a qubit $v$ can be regarded as removing the qubit after a process of local complementation (LC) on $v$, denoted as $\tau_{v}(G)$. Specifically, LC is defined as: among all neighbors of qubit $v$ in its resource state, if there was an edge between a pair of neighbors, then that edge is deleted; otherwise, an edge is added between that pair of neighbors.
To remove these cyclic structures, resource states with failed root-leaf fusions can be transformed to their local complementations of star-like structures (from $B$ to $C$ in Fig. 8) by applying the following sequence of 1-qubit operators (LCRule, )
$U_{v}(G)=\exp\large(-\mathrm{i}\frac{\pi}{4}X_{v}\large)\prod_{u\in N_{v}}\exp\large(\mathrm{i}\frac{\pi}{4}Z_{u}\large)$
with $N_{v}$ being the neighbors of $v$ in $G$. Computation on these local complementation states is equivalent to that on the original states. This is because we can interchange the orders of measurements and LC operators by adjusting the measurement bases, with the rules summarized in Theorem. 4.1. Similarly, the LC operators can also be interchanged with fusion operations, with the rules summarized in Theorem 4.2. Consequently, all LC operators can be postponed to the end of the computing process, which eliminates the necessity to implement them in the real-time.
@@
Theorem 4.1.
The local operator $U_{Z}^{\pm}=\exp(\pm\mathrm{i}\frac{\pi}{4}Z)$ or $U_{X}^{\pm}=\exp(\pm\mathrm{i}\frac{\pi}{4}X)$ can be propagated through a $Z-$measurement or a 1-qubit equatorial measurement on the Bloch sphere, i.e., a measurement in the basis of $\cos\phi X+\sin\phi Y$ where $\phi\in[0,2\pi)$, by a change of measurement basis.
@@
When measuring a 1-qubit state $\left|\psi\right\rangle$ along $A$-basis, the state collapses to $\left|\psi^{\prime}\right\rangle\equiv\mathrm{M}_{[A]}\left|\psi\right\rangle=\frac{\mathbb{I}\pm A}{2}\left|\psi\right\rangle$, with the sign $\pm$ determined by the measurement outcome, 0 or 1. Therefore,
$\mathrm{M}_{Z}U_{Z}^{\pm}$ $=U_{Z}^{\pm}\mathrm{M}_{Z}$
$\mathrm{M}_{Z}U_{X}^{\pm}$ $=U_{X}^{\pm}\mathrm{M}_{[\mp Y]}$
@@
$\mathrm{M}_{[\cos\phi X+\sin\phi Y]}U_{X}^{\pm}$ $=U_{X}^{\pm}\mathrm{M}_{[\cos\phi X\pm\sin\phi Z]}\qed$
Theorem 4.2.
The local operator $U_{Z}^{\pm}=\exp(\pm\mathrm{i}\frac{\pi}{4}Z)$ or $U_{X}^{\pm}=\exp(\pm\mathrm{i}\frac{\pi}{4}X)$ can be propagated through a 2-qubit $XZ,ZX$ fusion, i.e., a joint measurement of $X_{1}Z_{2},Z_{1}X_{2}$ on qubit 1 and qubit 2, by a change of fusion basis.
@@
When measuring a 2-qubit state $\left|\psi\right\rangle$ along basis $A_{1}B_{2}$, the state collapses to $\left|\psi^{\prime}\right\rangle\equiv\mathrm{M}_{[A_{1}B_{2}]}\left|\psi\right\rangle=\frac{\mathbb{I}\pm A_{1}B_{2}}{2}\left|\psi\right\rangle$, with $\pm$ determined by the measurement outcome, 0 or 1. Therefore
@@
$\mathrm{M}_{[X_{1}Z_{2}]}$ $\mathrm{M}_{[Z_{1}X_{2}]}U_{X_{1}}^{\pm_{1}}U_{X_{2}}^{\pm_{2}}=U_{X_{1}}^{\pm_{1}}U_{X_{2}}^{\pm_{2}}\mathrm{M}_{[\mp_{1}Y_{1}X_{2}]}\mathrm{M}_{[\mp_{2}X_{1}Y_{2}]}$
@@
4.3. Collective Feed-forward
The semi-static fusion strategy allows collective feed-forward in the granularity of RSL, which can be pipelined to reduce the overhead. On one hand, the propagation of LC operators through measurements and fusions requires an adaptive adjustment of measurement and fusion bases. With this dependency, each RSL are fused in two batches: a batch of root-leaf fusions and a batch of leaf-leaf fusions. On the other hand, the connectivity of the physical graph state can be enhanced by retries of the failed connections, including retrying failed leaf-leaf fusions with redundant degrees (e.g., with the $7^{\mathrm{th}}$ degree of $A_{1}$ in Fig. 8) and retrying failed root-leaf fusions with remaining degrees (e.g., with $A$ and $C$ in Fig. 8). In this way, each RSL may undergo more batches of fusions. However, since earlier batches of later RSLs can be conducted concurrently with later batches of earlier RSLs, this only introduces a constant overhead to program execution.
@@ 5. Random State Reshaping
Figure 9. (2+1)-D reshaping for handling random graph states generated by fusions.
In this section, we delve into the reshaping of physical graph states, which is characterized by a (2+1)-D design, motivated by the continuous generation of RSLs over time and the presence of delay lines. In Fig. 2, this corresponds to the online pass from physical graph states to measurement patterns.
@@ 5.1. Efficient 2D Renormalization
On each (merged) RSL, we apply a process known as renormalization (2D_renorm, ), which reshapes the largest connected component of the physical graph state to a coarse-grained 2D lattice. The key to its viability lies in the percolation phenomenon (3GHZ, ; perc_threshold, ; 2D_renorm, ). That is, when the fusion success probability exceeds a certain threshold, the random physical graph state undergoes a phase transition from short-range connectivity to long-range connectivity, leading to the largest connected component reaching a comparable size with the original graph state. Since fusions on each RSL are constrained as a squared lattice, the percolation threshold is only 0.5 (square_lattice_perc, ), lower than the achievable fusion success probability.
Identifying intersections of horizontal and vertical paths in the largest connected component reveals a coarse-grained square lattice, represented by bold nodes and edges in Fig. 9(b). This is achieved in the following way. We search for vertical paths from left to right and horizontal paths from bottom to top, enforcing distinct vertical or horizontal paths to maintain a separation of at least one qubit. When searching for vertical (horizontal) paths, a connectivity check is conducted between nodes at the top (left) and bottom (right), facilitated by a disjoint-set data structure to reduce the complexity. Upon confirming connectivity, a breadth-first search (BFS) is applied to determine the shortest path, ensuring it remains free of self-tangling. To further prevent tangling between vertical and horizontal paths, we remove the surrounding qubits of each identified path after discovery, preventing their interference with subsequent searches. Considering the removals, an alternating search of vertical and horizontal paths emerges as an effective searching order.
To improve real-time scalability, the 2D renormalization is designed to allow modularity, with areas on the RSL renormalized concurrently and then joined together. As shown in Fig. 10, the RSL is divided into several modules of size $L_{Module}\times L_{Module}$, with some intervals of length $L_{interval}$ left in between for joining the modules by connected paths. With the path searching algorithm above, the complexity of a modular 2D renormalization is $O(L_{\mathrm{module}}^{2})\sim O(N^{2}/m)$, where $m$ is the number of modules. Since an entire path can only be established if all inter-module paths involved are successful (e.g., the orange path), the potential for failed inter-module paths could lead to the renormalized lattice size being smaller than the total size of all individual modules. A suitable ratio of $L_{Module}$ and $L_{Interval}$, defined as MI ratio, can help mitigate this resource overhead.
@@
5.2. Flexible Time-like Connections
Nodes on the renormalized 2D lattices can be connected along the time dimension, referred to as time-like connections. Connections between adjacent 2D lattices and across non-adjacent 2D lattices are called adjacent-layer connections and cross-layer connections, respectively. Before program execution, the connections to establish are given by the IR program as a 3D graph, as illustrated in Fig. 9(a).
The process of generating this 3D graph is illustrated in Fig. 9(c) on 8 RSLs, with the adjacent-layer connection $AC$ and the cross-layer connection $BD$ implemented through the bold blue paths. This involves an attempt of 2D renormalization on each RSL. The successful ones then serve as logical layers, indexed by integers in Fig. 9(c), with the renormalized nodes on them referred to as logical nodes. In contrast, the RSLs with failed renormalization serve as routing layers, which are indexed by decimals in Fig. 9(c). The renormalization on an RSL is considered successful if:
@@ 5.2. Flexible Time-like Connections
    The renormalized 2D lattice reaches a target size, which is equivalent to a choice of average node size, where
    $\mathrm{average\_node\_size}\equiv\frac{\mathrm{RSL\_size}}{\mathrm{renormalized\_lattice\_size}}$
@@
    The RSL can establish all necessary time-like connections with prior logical layers through the following procedure.
@@
To establish a time-like connection between two nodes, a set of physical qubits around the preceding node are fused with corresponding qubits on a correct subsequent RSL. For adjacent-layer connections such as $AC$, the qubits around $A$ are directly fused to the next RSL, which is layer 1.1 in Fig. 9(c). For cross-layer connections such as $BD$, the qubits around $B$ are temporarily stored in delay lines, depicted by the green thin lines in Fig. 9(c), until they can be fused to layer 2.1, which is the first RSL between the current attempting RSL (layer 3) and its prior logical layer (layer 2). Subsequently, a path searching between the two nodes is conducted within the physical graph state, exemplified by the bold blue lines $AC$ and $BD$ in Fig. 9(c). Again, this is achieved by a connectivity check utilizing a disjoint-set data structure and a BFS for the shortest path. If the connectivity check yields a negative result, it indicates that the current RSL fails to meet the second condition and would become a routing layer. It is worth mentioning that the reshaping process can tolerate photon loss, since a fusion is considered as successful only if both two photons are detected. Effectively, the presence of photon loss causes a reduction of the fusion success probability, possibly leading to more routing layers between logical layers.
In contrast to the logical layers, all qubits of each routing layer are directly fused with their next RSL, as depicted by the grey thin lines in Fig. 9(c). This is because before obtaining the next successful renormalization, we can’t predict where the logical node would locate and which fusions around it would succeed. Moreover, in contrast to the simple case in Fig. 9 where there is only one connection between layer 1 and layer 2, in practice we may need to establish multiple connections between logical layers. This makes it even harder to predict which fusions are redundant before executing the fusions.
6. Offline Optimization with IR
In this section, we introduce the virtual hardware, FlexLattice IR, offline mapping and intermediate-level instructions. This covers the offline pass in the compilation flow (Fig. 2).
Before program execution, the mismatch between program graph states and physical graph states can be addressed by mapping the program graph state onto the virtual hardware, leading to an IR graph state that maintains the high-level program information. This IR program can then be transformed to a set of intermediate-level instructions, which guides real-time physical operations through the reshaping algorithm above.
6.1. Virtual Hardware
The virtual hardware abstracts the adjustable structures supported by the reshaping algorithm. It pocesses a (2+1)-D structure, characterized by the following features, as illustrated in Fig. 11(b).
  1. (1)
    The virtual hardware consists of consecutive layers of 2D lattices in a fixed size, with a virtual memory located on each 2D coordinate.
  2. (2)
    Nodes on the same 2D coordinate of different layers, either on adjacent or non-adjacent layers, can be connected along the third dimension, with the connections between non-adjacent layers realized by temporary storage of nodes in the virtual memory.
  3. (3)
    Each connection within or between 2D layers can be enabled or disabled on demand, but each node can have at most one connection with preceding layers and at most one connection with subsequent layers.
While this virtual hardware can be used to generate 3D cluster states (i.e., lattice-like graph states), which serve as the universal computing resource of MBQC in previous work (mbqc2009, ), it is more advantageous in the following aspects. First, an individual connection can be flexibly enabled or disabled without removing any logical node or affecting other edges. This is in contrast to the cluster state, wherein the removal of edges is usually achieved by removing involved vertices and all their edges. Second, the connections among 2D layers exhibit greater flexibility than cluster states. Specifically, inter-layer connections between nodes on the same 2D coordinates extend beyond adjacent layers, encompassing cross-layer connections as well.